# On Generalized Characters Whose Values on Nonidentity Elements Are Sums of at most Two Roots of Unity

### v2

@@ 1. Introduction
Theorem 2 is also applicable to certain classes of permutation groups. A permutation group $G$ acting on a set $\Omega$ is said to have minimal degree $f$ if each $g\in G^{\#}$ fixes at most $|\Omega|-f$ points in $\Omega$. The transitive permutation groups of minimal degree $|\Omega|$ are precisely those which act regularly. Those of minimal degree $|\Omega|-1$. include the class of Frobenius groups. Those of minimal degree $|\Omega|-2$ include the class of Zassenhaus groups, i.e., doubly transitive groups having no nonidentity element fixing three points simultaneously. These Zassenhaus groups include the groups $\operatorname{PSL}(2,q)$ for $q$ odd along with the Suzuki groups, $\operatorname{Sz}(q)$.
@@
We do not get a full converse to this theorem. Specifically, let $G$ be the semidirect product of the cyclic group of order 15, $C_{15}$, with the cyclic group of order 2, $C_{2}$, where the involution acts as the automorphism of inversion. The prime graph has two connected components since there exists a subgroup of order 15 and the elements of order two do not commute with the subgroup of order 15. Let $1=\lambda_{1},\lambda_{2}\in\operatorname{Irr}(C_{15})$ be the two nontrivial irreducible characters of $C_{15}$ whose kernels contain the cyclic group of order 5. Define $\chi(x)=\lambda_{1}(x)+\lambda_{2}(x)-1$ for each $x\in C_{15}^{\#}$. We also take $\chi(x)=0$ for each involution $x$. Lastly, define $\chi(1)=16$. Brauer’s Characterization of Characters implies that $\chi$ is a generalized character. This character is a sum of at most two roots of unity on each nonidentity element. In particular, there exists $x,y,z\in G^{\#}$ such that $\chi(x)=0$, $\chi(y)=1$, and $\chi(z)=-2$. Yet the prime graph has only two connected components: $\{2\}$ and $\{3,5\}$.
@@ 2. Preliminaries
    If $\chi$ is a root of unity for all nonidentity $p$-elements, then when $p$ is odd or $|P|=2$, $\chi(1)\equiv\pm 1\ (\mathrm{mod}\ |P|)$. Otherwise, $p=2$ and $|P|\geq 4$, and $\chi(1)\equiv\pm 1\ (\mathrm{mod}\ |P|/2)$.
@@
By adding integer multiples of the regular character to $\chi$, we ma assume that $\chi$ is in fact a character of $G$ without altering our assumptions. Thus, we will assume $\chi$ is a character below.
@@
Assume that $\lambda_{1}$ takes at last 3 distinct values. By multiplying each character by $\overline{\lambda}_{3}$, we can assume without loss that $\lambda_{3}$ is trivial. It now suffices to produce an element $g\in G-(\ker\lambda_{1}\cup\ker\lambda_{2})$ (which is guaranteed to be nonempty) for which $\lambda_{1}(g)\neq\lambda_{2}(g)$. If $\ker\lambda_{1}\subseteq\ker\lambda_{2}$ or vice versa, the statement follows by the distinctness of the two characters, so assume otherwise. In particular, we may assume that neither of the two characters are faithful.
@@ 3. Proofs of Main Result
Assume that $\chi$ is a counterexample to our main theorem. We may subtract rational (or even real) multiples of the regular character from $\chi$ without affecting the values $\chi$ takes on $G^{\#}$. Thus, we may assume that $\chi(1)=0$. Now, we may write
@@
Assume there exists $\eta,\theta\in\operatorname{Irr}(G)$ such that $|a_{\eta}-a_{\theta}|\geq 3$. We have that
$a_{\eta}^{2}+(a_{\eta}\pm 3)^{2}\leq a_{\eta}^{2}+a_{\theta}^{2}\leq[\chi,\chi]<4.$
However, by treating $a_{\eta}^{2}+(a_{\eta}\pm 3)^{2}$ as a polynomial in the variable $a_{\eta}$, one sees that the minimum value taken by this expression is $9/2$, contradicting the above inequality. Thus, there exists $x\in\mathbb{Q}$ such that $\chi,\lambda]=x+i$, for $i=0,1,2$ and for each $\lambda\in\operatorname{Irr}(G)$.
@@
A valid choice of $x$ will not exist if this discriminant is negative, so we assume this to be the case. Algebraically rearranging, we obtain
$4<\dfrac{a_{0}a_{1}+a_{1}a_{2}+4a_{0}a_{2}}{a_{0}+a_{1}+a_{2}-1}.$
If either $a_{0}$ or $a_{2}$ equals one, then one verifies that this inequality forces the other to be at most 3. Notice that if $a_{0}=1$ and $a_{2}=3$ (or vice versa), equality actually holds. Likewise, if $a_{0}$ and $a_{2}$ are both at least 2, then the above inequality holds, so this case cannot hold either. On the other hand, if $a_{2}=0$, then the above inequality becomes
(3) $4(a_{0}+a_{1}-1)-a_{0}a_{1}<0.$
@@
In this case, $\chi=a\rho\pm(\lambda_{1}+\lambda_{2}+\lambda_{3})$ for distinct $\lambda_{i}\in\operatorname{Irr}(G)$. Since the sign of $a$ is irrelevant, it is of no loss to multiply $\chi$ by a sign, so we may assume $\chi=a\rho+\lambda_{1}+\lambda_{2}+\lambda_{3}.$ Now, choose $g\in G$. We have that
@@
Lastly, if at exactly one of $\delta_{1}=\pm 1$ and $\delta_{2}=0$ (or vice versa), then we have $\lambda_{1}(g)+\lambda_{2}(g)+\lambda_{3}(g)\pm\epsilon_{1}=0$, a vanishing sum of weight 4. This can only be partitioned into two vanishing sums of weight 2, and in this case, at least two of the terms will have even order. If $o(g)$ is odd, then each $\lambda_{i}$ will have odd order, an impossibility. Thus, $o(g)$ is even, and $2\in\pi(G)$.
@@
Again, we may assume without loss that $\chi=a\rho+\lambda_{1}+\lambda_{2}-\lambda_{3}$ for distinct $\lambda_{i}\in\operatorname{Irr}(G)$. Again, choose $g\in G$. We may write
@@
We have seen that $a_{0}\geq 5$ implies $a_{2}=0$ and vice versa. Thus, we may assume without loss that $a_{2}=0$ and $a_{1}\geq 5$. We have seen that this implies $|G|\leq 21$. The only groups of order at most 21 and indivisible by either 2, 3, or 5 are the cyclic groups of order 7, 11, 13, 17, and 19. Further analysis of the inequality in (3) shows that $a_{1}=6$ implies $a_{0}\leq 10$ and that $a_{1}=7$ implies $a_{0}\leq 8$. Cases with higher values of $a_{1}$ are subsumed by the cases considered here by symmetry. In particular, it follows that in this case, $\chi=a\rho_{G}\pm\sum_{i=1}^{k}\lambda_{i}$ for distinct $\lambda_{i}\in\operatorname{Irr}(G)$ and $5\leq k\leq 7$.
If $a_{1}=5$, then considering $\chi(g)$ for any element $g\in G^{\#}$ gives rise to a vanishing sum of weight between 5 and 7. Since at least five of the terms will have odd order, it follows that that $G$ is the cyclic group of order 7, but in this case, $\chi$ is some multiple of the regular character minus two linear characters. Similarly, it follows that $|G|=7$ when considering the cases where $a_{1}$ is 6 or 7, meaning that $\chi$ is not a meaningful counterexample. This completes the proof.
@@ 4. Prime Graphs
Now, assume there exists an abelian subgroup $H$ of order $pq$ such that $\chi_{H}$ is not of the form $a\rho_{H}+\delta_{1}\lambda_{1}+\delta_{2}\lambda_{2}$ for $\lambda_{i}\in\operatorname{Irr}(H)$, $a\in\mathbb{Z}$, and $\delta_{i}\in\{-1,0,1\}$. In particular, our main result implies that $p=5$. Further , since $|H|=5q\geq 35$, $\chi_{H}$ is of the form of one of the four outlier cases. Set $H_{5}$ and $H_{q}$ as the Sylow $5$- and Sylow $q$-subgroups respectively. We proceed by cases.
@@
In the former cases, as in the first paragraph of this proof, we obtain that $\chi(1)\equiv\pm 1\ (\mathrm{mod}\ 5)$ and (modq)\ ($\mathrm${mod}\ q). That is, either $p,q\in\pi_{1}$ or $p,q\in\pi_{-1}$. Likewise, in the latter cases, we have $\chi(1)\equiv\pm 2\ (\mathrm{mod}\ 5)$ and (modq)\ ($\mathrm${mod}\ q), and $p,q\in\pi_{2}$ or $p,q\in\pi_{-2}$.
We now obtain a result about the connectivity of the induced subgraph of $\Gamma(G)$ on the vertex set $\pi(G)-\{2\}$. First, we introduce some notation. Let $H$ be an abelian group, and let $\chi$ be a generalized character of $H$ taking the form
@@
We at last have our final result on prime graphs which says that 2 is often a “cut vertex” of the prime graph under our assumptions, that is, the prime graph becomes disconnected when discarding the vertex 2.
@@
We proceed in cases depending on the congruence class of $\chi(1)$ modulo 3 and modulo 5. Using the notation of the previous theorem, we will take repeated advantage of the fact that the induced subgraph of $\Gamma^{\prime\prime}(G)$ on the vertex set $\pi_{i}$ is either empty or a union of connected components of $\Gamma^{\prime\prime}(G)$.
@@
However, if $\chi_{H}$ has the latter two types, this implies $|H|\leq 16$ by Corollary 13, and so $p=5$, but this is impossible by the above lemma. If $\chi_{H}$ has either of the first two types, this implies that 3 is linked to an element of $\pi_{0}$. If $G$ has another subgroup $K$ of order $3q$ for some prime $q$ such that $\chi_{K}$ has type $(3,0)$ or $(0,3)$, then $p=5$ by our previous lemma. However, these two types cannot occur within $G$ simultaneously since this would imply that $\chi(1)\equiv-2\ (\mathrm{mod}\ 5)$ and $\chi(1)\equiv 2\ (\mathrm{mod}\ 5)$. In any case, 3 can at most link $\pi_{0}$ to either $\pi_{-2}$ or $\pi_{2}$, but not both simultaneously. More importantly, there is no path in $\Gamma^{\prime}(G)$ linking an element of $\pi_{1}\cup\pi_{-1}$ to an element of $\pi_{0}\cup\pi_{2}\cup\pi_{-2}$.
Now, assume that $\chi(1)\equiv 1\ (\mathrm{mod}\ 3)$. If $H$ is an abelian subgroup of order $3p$, then $\chi_{H}$ can only have the following types:
@@
Let $K$ be an abelian subgroup of type $3q$ for some prime $q$. If $\chi_{H}$ and $\chi_{K}$ have types $(1,0)$ and $(0,2)$ respectively, then this implies that there possibly exists a path between an element of $\pi_{1}$ and an element of $\pi_{-2}$ passing through 3. Our previous results imply that these are the only possible cases to consider if $p$ and $q$ are greater than 7. Also, if $G$ has a cyclic subgroup $C$ of order $21$ such that $\chi_{C}$ has type $(0,5)$, then this implies $\chi(1)\equiv 2\ (\mathrm{mod}\ 7)$, and so, 3 may be connected by a path to $\pi_{2}$.
Assume that $G$ has a cyclic subgroup $C$ of order 15. We proceed by cases depending on the congruence class of $\chi(1)$ modulo $5$. If $5\mid\chi(1)$, this implies that $\chi_{C}$ can only have type $(0,5)$. However, this is impossible by our previous lemma. If $\chi(1)\equiv-2\ (\mathrm{mod}\ 5)$, then the only possible exceptional type for $\chi_{C}$ is $(1,3)$, but this implies $5\in\pi_{-2}$. Likewise, if $\chi(1)\equiv 1\ (\mathrm{mod}\ 5)$, then the only possible exceptional type for $\chi_{C}$ is $(2,1)$, but this implies that $5\in\pi_{1}$. The existence of paths from 3 to elements of these sets was already established in the previous paragraph. If $\chi(1)\equiv 2\ (\mathrm{mod}\ 5)$, then the only exceptional type possible for $\chi_{C}$ is $(7,0)$. This implies the existence of a possible path connecting 3 to an element of $\pi_{2}$, but again, this was already established in the previous paragraph. Lastly, if $\chi(1)\equiv-1\ (\mathrm{mod}\ 5)$, then the only exceptional type for $\chi_{C}$ is $(4,0)$ implying the potential existence of a connection between 3 and $\pi_{-1}$. In any case where $\chi(1)\equiv 1\ (\mathrm{mod}\ 3)$, there is no path connecting 3 to an element of $\pi_{0}$. Thus, $\Gamma^{\prime}(G)$ is disconnected in this case as well.

### v3

@@ 1. Introduction
Theorem 2 is also applicable to certain classes of permutation groups. A permutation group $G$ acting on a set $\Omega$ is said to have minimal degree $f$ if each $g\in G^{\#}$ fixes at most $|\Omega|-f$ points in $\Omega$. The transitive permutation groups of minimal degree $|\Omega|$ are precisely those which act regularly. Those of minimal degree $|\Omega|-1$ include the class of Frobenius groups. Those of minimal degree $|\Omega|-2$ include the class of Zassenhaus groups, i.e., doubly transitive groups having no nonidentity element fixing three points simultaneously. These Zassenhaus groups include the groups $\operatorname{PSL}(2,q)$ for $q$ odd along with the Suzuki groups, $\operatorname{Sz}(q)$.
@@
We do not get a full converse to this theorem. Specifically, let $G$ be the semidirect product of the cyclic group of order 15, $C_{15}$, with the cyclic group of order 2, $C_{2}$, where the involution acts as the automorphism of inversion. The prime graph has two connected components since there exists a subgroup of order 15 and the elements of order two do not commute with the subgroup of order 15. Let $\lambda_{1},\lambda_{2}\in\operatorname{Irr}(C_{15})$ be the two nontrivial irreducible characters of $C_{15}$ whose kernels contain the cyclic group of order 5. Define $\chi(x)=\lambda_{1}(x)+\lambda_{2}(x)-1$ for each $x\in C_{15}^{\#}$. We also take $\chi(x)=0$ for each involution $x$. Lastly, define $\chi(1)=16$. Brauer’s Characterization of Characters implies that $\chi$ is a generalized character. This character is a sum of at most two roots of unity on each nonidentity element. In particular, there exists $x,y,z\in G^{\#}$ such that $\chi(x)=0$, $\chi(y)=1$, and $\chi(z)=-2$. Yet the prime graph has only two connected components: $\{2\}$ and $\{3,5\}$.
@@ 2. Preliminaries
    If $\chi$ is a root of unity for all nonidentity $p$-elements, then when $p$ is odd or $|P|=2$, $\chi(1)\equiv\pm 1\ (\mathrm{mod}\ |P|)$. Otherwise, $p=2$, $|P|\geq 4$, and $\chi(1)\equiv\pm 1\ (\mathrm{mod}\ |P|/2)$.
@@
By adding integer multiples of the regular character to $\chi$, we may assume that $\chi$ is in fact a character of $G$ without altering our assumptions. Thus, we will assume $\chi$ is a character below.
@@
By multiplying each character by $\overline{\lambda}_{3}$, we can assume without loss that $\lambda_{3}$ is trivial. It now suffices to produce an element $g\in G-(\ker\lambda_{1}\cup\ker\lambda_{2})$ (which is guaranteed to be nonempty) for which $\lambda_{1}(g)\neq\lambda_{2}(g)$. If $\ker\lambda_{1}\subseteq\ker\lambda_{2}$ or vice versa, the statement follows by the distinctness of the two characters, so assume otherwise. In particular, we may assume that neither of the two characters are faithful.
@@ 3. Proofs of Main Result
Assume that $\chi\in\mathbb{Z}[\operatorname{Irr}(G)]$ satisfies the conditions of the main theorem. We may subtract rational (or even real) multiples of the regular character from $\chi$ without affecting the values $\chi$ takes on $G^{\#}$. Thus, we may assume that $\chi(1)=0$. Now, we may write
@@
Assume there exists $\eta,\theta\in\operatorname{Irr}(G)$ such that $a_{\eta}-a_{\theta}\geq 3$. We have that
$a_{\theta}^{2}+(a_{\theta}\pm 3)^{2}\leq a_{\theta}^{2}+a_{\eta}^{2}\leq[\chi,\chi]<4.$
However, by treating $a_{\theta}^{2}+(a_{\theta}\pm 3)^{2}$ as a polynomial in the variable $a_{\theta}$, one sees that the minimum value taken by this expression is $9/2$, contradicting the above inequality. Thus, there exists $x\in\mathbb{Q}$ such that $\chi,\lambda]=x+i$, for $i=0,1,2$ and for each $\lambda\in\operatorname{Irr}(G)$.
@@
A valid choice of $x$ will not exist if this discriminant is negative, so we assume this not to be the case. Algebraically rearranging, we obtain
$4\geq\dfrac{a_{0}a_{1}+a_{1}a_{2}+4a_{0}a_{2}}{a_{0}+a_{1}+a_{2}-1}.$
If either $a_{0}$ or $a_{2}$ equals one, then one verifies that this inequality forces the other to be at most 3. Notice that if $a_{0}=1$ and $a_{2}=3$ (or vice versa), equality actually holds. Likewise, if $a_{0}$ and $a_{2}$ are both at least 2, then the above inequality fails, so this case cannot hold either. On the other hand, if $a_{2}=0$, then the above inequality becomes
(3) $4(a_{0}+a_{1}-1)-a_{0}a_{1}\geq 0.$
@@
In this case, $\chi=a\rho\pm(\lambda_{1}+\lambda_{2}+\lambda_{3})$ for distinct $\lambda_{i}\in\operatorname{Irr}(G)$. Since the sign of $a$ is irrelevant, it is of no loss to multiply $\chi$ by a sign, so we may assume $\chi=a\rho+\lambda_{1}+\lambda_{2}+\lambda_{3}.$ Now, choose $g\in G^{\#}$. We have that
@@
Lastly, if at exactly one of $\delta_{1}=\pm 1$ and $\delta_{2}=0$ (or vice versa), then we have $\lambda_{1}(g)+\lambda_{2}(g)+\lambda_{3}(g)\pm\epsilon_{1}=0$, a vanishing sum of weight 4. This can only be partitioned into two vanishing sums of weight 2 by Lemma 11, and in this case, at least two of the terms will have even order. If $o(g)$ is odd, then each $\lambda_{i}$ will have odd order, an impossibility. Thus, $o(g)$ is even, and $2\in\pi(G)$.
@@
Again, we may assume without loss that $\chi=a\rho+\lambda_{1}+\lambda_{2}-\lambda_{3}$ for distinct $\lambda_{i}\in\operatorname{Irr}(G)$. Again, choose $g\in G^{\#}$. We may write
@@
We have seen that $a_{0}\geq 5$ implies $a_{2}=0$ and vice versa. Thus, we may assume without loss that $a_{2}=0$ and $a_{1}\geq 5$. We have seen that this implies $|G|\leq 21$. The only groups of order that are at most 21 and are indivisible by either 2, 3, or 5 are the cyclic groups of order 7, 11, 13, 17, and 19. Further analysis of the inequality in (3) shows that $a_{1}=6$ implies $a_{0}\leq 10$ and that $a_{1}=7$ implies $a_{0}\leq 8$. Cases with higher values of $a_{1}$ are subsumed by the cases considered here by symmetry. In particular, it follows that in this case, $\chi=a\rho_{G}\pm\sum_{i=1}^{k}\lambda_{i}$ for distinct $\lambda_{i}\in\operatorname{Irr}(G)$ and $5\leq k\leq 7$.
If $a_{1}=5$, then considering $\chi(g)$ for any element $g\in G^{\#}$ gives rise to a vanishing sum of weight between 5 and 7. Since over half of the terms will have odd order, it follows by Lemma 9 that the terms of $\chi(g)$ are 7-th roots of unity, and so $|G|=7$. In this case however, $\chi$ is some multiple of the regular character minus two linear characters. Similarly, it follows that $|G|=7$ when considering the cases where $a_{1}$ is 6 or 7, meaning that $\chi$ is not a meaningful counterexample. This completes the proof.
@@ 4. Prime Graphs
Now, assume there exists an abelian subgroup $H$ of order $pq$ such that $\chi_{H}$ is not of the form $a\rho_{H}+\delta_{1}\lambda_{1}+\delta_{2}\lambda_{2}$ for $\lambda_{i}\in\operatorname{Irr}(H)$, $a\in\mathbb{Z}$, and $\delta_{i}\in\{-1,0,1\}$. In particular, our main result implies that $p=5$. Further , since $|H|=5q>21$, $\chi_{H}$ is of the form of one of the four outlier cases. Set $H_{5}$ and $H_{q}$ as the Sylow $5$- and Sylow $q$-subgroups respectively. We proceed by cases.
@@
In the former cases, as in the first paragraph of this proof, we obtain that $\chi(1)\equiv\pm 1\ (\mathrm{mod}\ 5)$ and (modq)\ ($\mathrm${mod}\ q). That is, either $p,q\in\pi_{1}$ or $p,q\in\pi_{-1}$. Likewise, in the latter cases, we have $\chi(1)\equiv\pm 2\ (\mathrm{mod}\ 5)$ and (modq)\ ($\mathrm${mod}\ q), and so, either $p,q\in\pi_{2}$ or $p,q\in\pi_{-2}$.
We now may obtain a result on the connectivity of the induced subgraph of $\Gamma(G)$ on the vertex set $\pi(G)-\{2\}$. First, we introduce some notation. Let $H$ be an abelian group, and let $\chi$ be a generalized character of $H$ taking the form
@@
We at last have our final result.
@@
Using the notation of the previous theorem, we take advantage of the fact that the induced subgraph of $\Gamma^{\prime\prime}(G)$ on the vertex set $\pi_{i}$ is either empty or a union of connected components of $\Gamma^{\prime\prime}(G)$. Thus, if any element of $\pi_{i}$ becomes connected by a path to an element of $\pi_{j}$ for $i\neq j$, this path must pass through the vertex 3. We proceed in cases depending on the congruence class of $\chi(1)$ modulo 3. In each case, we investigate the possible primes $p$ such that a cyclic subgroup of order $3p$ can exist within $G$.
@@
However, if $\chi_{H}$ has the latter two types, this implies $|H|\leq 16$ by Corollary 13, and so $p=5$, but this is impossible by the above lemma. If $\chi_{H}$ has either of the first two types, this implies that 3 is linked to an element of $\pi_{0}$. If $G$ has another subgroup $K$ of order $3q$ for some prime $q$ such that $\chi_{K}$ has type $(3,0)$ or $(0,3)$, then $p=5$ by our previous lemma. However, these two types cannot occur within $G$ simultaneously since this would imply that $\chi(1)\equiv-2\ (\mathrm{mod}\ 5)$ and $\chi(1)\equiv 2\ (\mathrm{mod}\ 5)$. In any case, 3 can at most link $\pi_{0}$ to either $\pi_{-2}$ or $\pi_{2}$, but not both simultaneously. More importantly, there is no path in $\Gamma^{\prime}(G)$ linking an element of $\pi_{1}\cup\pi_{-1}$ to an element of $\pi_{0}\cup\pi_{2}\cup\pi_{-2}$, and the result follows in this case.
Now, assume that $\chi(1)\equiv 1\ (\mathrm{mod}\ 3)$. We claim that the induced subgraph of $\Gamma(G)$ on $\pi_{0}$ is a union of connected components of $\Gamma^{\prime}(G)$. By the previous theorem, it suffices to show that there exists no cyclic subgroup of order $3p$ for $p\in\pi_{0}$. If $H$ is an abelian subgroup of order $3p$, then $\chi_{H}$ can only have the following types:
@@
If $\chi_{H}$ has type $(k,\ell)$ and $p\in\pi_{0}$, then $\chi(1)=a|H|+k$-$\ell$ is divisible by $p$, but since $p\mid|H|$, we have that $p\mid(k-\ell)$. Given the possible types $(k,\ell)$ above, the only prime divisors greater than 3 for any choice of $k$-$\ell$ are $5$ and $7$. If $p=5$, then the only possible type for $\chi_{H}$ is $(0,5)$, but by Lemma 16, this is impossible since $|H|=15$ in this case. On the other hand, if $p=7$, then the only possible type for $\chi_{H}$ is $(7,0)$, but by Corollary 13, this type can only occur if $|H|\leq 15$ despite the fact that $|H|=21$ in this case. Thus, $3$ is not not linked to any element of $\pi_{0}$ in $\Gamma^{\prime}(G)$. However, by Theorem 15, there exist no edges between an element of $\pi_{0}$ and an element of $\pi_{i}$ for $i\neq 0$ in $\Gamma^{\prime}(G)$. Thus, the vertices of $\pi_{0}$ form a union of connected component of $\Gamma^{\prime}(G)$ in this case.